OpenStudy (anonymous):

how do you prove that eigenvalues and eigenvectors are correct for a given matrix?

OpenStudy (anonymous):

I have eigenvalues -1,1,2,4 with eigenvectors (0,0,1,0),(-2,2,1,0),(0,-2,-1,1), and (0,0,1,5) respectively

OpenStudy (anonymous):

the initial matrix is \[\left[\begin{matrix}1 & 0&0&0 \\ 1 & 2&0&0\\1&2&-1&1\\1&1&0&4\end{matrix}\right]\]

OpenStudy (anonymous):

is there any way to multiply them or something to obtain the initial matrix?

OpenStudy (turingtest):

you need to show that\[A\vec b=\lambda\vec b\]so once you have an eigenvector \(\vec\eta_n\) you want to show that after that vector undergoes the transformation due to \(A\) it will point in the same direction, i.e. be a constant multiple of itself.

OpenStudy (turingtest):

so\[A\vec\eta_i=\lambda_i\eta_i\]notice the change in magnitude is by a factor \(\lambda_i\), which is the \(i^{th}\) eigenvalue

OpenStudy (turingtest):

\[A\vec\eta_i=\lambda_i\vec\eta_i\]

OpenStudy (anonymous):

so it'll be vector=vector, right?

OpenStudy (turingtest):

A* eigenvector = constant * eigenvector

OpenStudy (turingtest):

you are showing that multiplying the eigenvector by the matrix is the same as multiplying the eigenvector by its eigenvalue

OpenStudy (turingtest):

for instance if my eigenvalue \(\lambda\) is 2, and my eigenvector \(\vec\eta\) is (1,-1), then you should be able to show that \[A\vec\eta=A\left[\begin{matrix}1\\-1\end{matrix}\right]=\left[\begin{matrix}2\\-2\end{matrix}\right]=2\left[\begin{matrix}1\\-1\end{matrix}\right]=\lambda\vec\eta\]

OpenStudy (anonymous):

okay, so for the first one, I get A*(vector) = 0 but that doesn't sound right

OpenStudy (turingtest):

I'm not sure, haven't tried the problem yet

OpenStudy (anonymous):

or [0//0//0//0]

OpenStudy (turingtest):

what's the eigenvalue you are looking at?

OpenStudy (anonymous):

-1

OpenStudy (turingtest):

let me try, I haven't done this in a bit

OpenStudy (anonymous):

so does lambda in this one have to be an eigenvalue?

OpenStudy (turingtest):

not sure I get you, lambda is an eigenvalue, yes

OpenStudy (anonymous):

I just plugged the rest into my calculator quick and they were all multiples of the eigenvector, but they weren't by one of the eigenvalues I had

OpenStudy (anonymous):

I end up getting *0, *2, *3, and *5 for lambda for the four of them

OpenStudy (turingtest):

so the eigenvector you got corresponding to lambda=-1 is what?

OpenStudy (anonymous):

wait, was I solving for something else? I took A times a 4x1 matrix of the eigenvector and solved for lambda.

OpenStudy (turingtest):

yeah, but what did you get for that eigenvector

OpenStudy (anonymous):

I got 0 for lambda

OpenStudy (turingtest):

I think we are having a laps in communication. You got your eigenvectors as {-1,1,2,4}, correct? then you got the eigenvector for each eigenvalue. what eigenvector did you get for the eigenvalue -1 ?

OpenStudy (anonymous):

sorry, the vector for -1 is (0,0,1,0)

OpenStudy (turingtest):

good that is what I got. now multiply that by A

OpenStudy (anonymous):

I got 0

OpenStudy (turingtest):

I have a sneaking suspicion that you are multiplying your eigenvector \(\vec\eta\) by \(A-\lambda\) and not \(A\)

OpenStudy (turingtest):

\[\left[\begin{matrix}1 & 0&0&0 \\ 1 & 2&0&0\\1&2&-1&1\\1&1&0&4\end{matrix}\right]\left[\begin{matrix}0\\0\\1\\0\end{matrix}\right]=?\]

OpenStudy (anonymous):

[0,0,-1,0]?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

which is\[-1\cdot\left[\begin{matrix}0\\0\\1\\0\end{matrix}\right]=\lambda\vec\eta\]

OpenStudy (anonymous):

okay. Then that would be a multiple of -1, which is the eigenvalue. I wonder what I did wrong...

OpenStudy (turingtest):

well it's right now, you see that, correct?

OpenStudy (anonymous):

ya I think so

OpenStudy (turingtest):

\[A\vec\eta=\left[\begin{matrix}1 & 0&0&0 \\ 1 & 2&0&0\\1&2&-1&1\\1&1&0&4\end{matrix}\right]\left[\begin{matrix}0\\0\\1\\0\end{matrix}\right]=\left[\begin{matrix}0\\0\\-1\\0\end{matrix}\right]=-1\cdot\left[\begin{matrix}0\\0\\1\\0\end{matrix}\right]=\lambda\vec\eta~~~~~~~~~~~~\huge\checkmark\]

OpenStudy (turingtest):

my guess is that you multiplied by \(A-\lambda\) and got\[(A-\lambda I)\vec\eta=\left[\begin{matrix}2 & 0&0&0 \\ 1 & 3&0&0\\1&2&0&1\\1&1&0&5\end{matrix}\right]\left[\begin{matrix}0\\0\\1\\0\end{matrix}\right]=\left[\begin{matrix}0\\0\\0\\0\end{matrix}\right]\]

OpenStudy (anonymous):

yep. I didn't realize that for some reason. Thanks for the help. I think I got it from here

OpenStudy (turingtest):

great, glad I could help :)